Caso de factor común por agrupación de términos. Factorizar:
b. 6x – 6y – by + bx
c. 8m3 – 2m2 – 2mn + 8m2n

Answer by stanbon(75887)

b. 6x – 6y – by + bx
= 6(x-y)-b(y-x)
= 6(x-y)+b(x-y)
= (x-y)(6+b)
======================
c. 8m3 – 2m2 – 2mn + 8m2n
Rearrange:
8m^3+8m^2n - 2m^2-2mn
-----
= 8m^2(m+n)-2m(m+n)
-----
= (m+n)(8m^2-2m)
= 2(m+n)(4m^2-m)
=================
Cheers,
Stan H.
=================